Drive home from Dallas at 60 mph and returned home at speed of 45 mph. If the total driving time was 21 hours, how far is it from your home to Dallas? I have tried the D=r*t formula but I can't seem to plug it together correctly. Can you show me the correct equation by steps? Thank you
Answer by drk(1908)

This is an RTD problem. creating a table from above, we get
driving . . . . . rate . . . . . .time . . . . . distance
from D . . . . . .60 . . . . . . .T . . . . . . . . 60T
home . . . . . . .45 . . . . . .21 - T . . . . . 945 - 45T
total . . . . . . . . . . . . . . . . 21. . . . . . . . . . . . . .
It appears that the distances are equal, so we get
60T = 945 - 45T
and then
105T = 945
T = 9.
We need distances, so
60(9) = 540 miles
45 (21-9) = 45(12) = 540 miles.